# Coning off totally geodesic boundary components of a hyperbolic manifold

### v1

@@ 5. Proof of Theorem B
We first show that $B_{\varepsilon}(y)$ is convex in $M$. Let $z,w\in B_{\varepsilon}(y)$. By the triangle inequality, any shortest path $[z,w]$ stays inside $B_{3\varepsilon}(y)$. Since $\varepsilon<\frac{1}{3}(b^{\prime}-t(y))$, $B_{3\varepsilon}(y)$ is contained in $C_{f}(N)$. In particular any shortest path in $X$ joining $z$ to $w$ is also a shortest path in the warped product metric on $C_{f}(N)$ (as opposed to a path leaving $C_{f}(N)$). By Lemma 4.1, $C_{f}(N)$ is $\operatorname{CAT}(0)$. In particular, balls in $C_{f}(N)$ are convex, so the segment $[z,w]$ lies inside $B_{\varepsilon}(y)$ as desired. This argument also shows
$d_{X}(z,w)=d_{C_{f}(N)}(z,w).$
We now assume $z,w$ are in $Y\cap B_{\varepsilon}(y)$. Let $u$ be a point on $[z,w]$. We must show $u\in Y$. Because the warping function $f$ is increasing, we always have $t(u)\leq\max\{t(z),t(w)\}$. To show $u\in Y$, we must show either that $p(u)\in P_{t(u)}$ (if $t(u)>b$), or that $p(u)\in P_{b}$ (if $t(u)\leq b$). To simplify, we introduce the following notation
@@
We have $P_{t(u)}\supseteq P_{t(z)}\cup P_{t(w)}$ by hypothesis 2, and so $P^{*}_{t(u)}\supseteq P^{*}_{t(z)}\cup P^{*}_{t(w)}$. Notably this means that the endpoints of the geodesic $[p(z),p(w)]$ are in $P^{*}_{t(u)}$. Since $z$ and $w$ are in $B_{\varepsilon}(y)$,
@@
Because $\varepsilon<\frac{1}{2}L=\frac{1}{3}(t(y)-t_{0})$, the ball $B_{\varepsilon}(y)$ is disjoint from $[t_{0},L]\times N\subset C_{f}(N)$. Since the warping function $f$ is increasing and $B_{\varepsilon}(y)$ is entirely above $\{L\}\times N$,
($\spadesuit$) $d_{X}(z,w)\geq d_{\{L\}\times N}((L,p(z)),(L,p(w)))=f(L)d_{\partial M}(p(z),p(w)).$
@@
Without loss of generality suppose $t(w)<t(u)<t(z)$. By the argument in Case 1, $[p(z),p(w)]\subseteq P^{*}_{t(w)}$. For the sake of contradiction, suppose that $u\in[z,w]$ is not in $Y$. Re-choosing $w$ if necessary, we may assume $t(\cdot)$ is monotone on $[z,w]$. Possibly shortening the segment again, we may assume that $[z,w]$ meets $Y$ only in its endpoints. Thus, for $t\in(t(w),t(z))$, we must have ${P}^{*}_{t}\neq P^{*}_{t(w)}$. Hence $t(w)\geq b$, since $P^{*}_{t}=P_{b}$ for all $t\leq b$. This means that $[z,w]$ is actually a shortest path in $M$, contained in some $2\varepsilon$–ball around a point in $S$. Similarly, the points $z,w$ are points of $S$. Since $\varepsilon<\frac{1}{2}\mathrm{lcr}_{S}(S)$, all of $[z,w]$ is contained in $S$, and hence in $Y$. In particular, $u\in Y$, a contradiction. ∎
@@
Let $y$ be the cone point of $C_{f}(N)$ for a component $N$ of $\partial M$ so that $N\cap S\neq\emptyset$. Let $\varepsilon=\frac{1}{3}(b-t_{0}).$ For any $z,w\in B_{\varepsilon}(y)$, the triangle inequality shows that a shortest path lies entirely in $C_{f}(N)$, and
@@
Lemma 4.1 shows $C_{f}(N)$ is $\operatorname{CAT}(0)$, so in fact there is a unique shortest segment $[z,w]$, which lies in $B_{\varepsilon}(y)$.
We suppose now that $z,w$ lie in $Y\cap B_{\varepsilon}(y)$. If $[z,w]$ goes through the cone point it is clearly contained in $Y$, so we assume it avoids the cone point. We can thus write $[z,w]$ as a path
@@
where the path $\rho(s)$ is a (not necessarily constant speed) geodesic in $\partial M$. We must show that $\rho(s)\in P_{b}$ for all $s$. By way of contradiction suppose some $\rho(s)\not\in P_{b}$. Replacing $[z,w]$ by a sub-segment if necessary, we may assume that $\rho(s)\not\in P_{b}$ for all $s\in(0,1)$. We therefore have
$d_{\partial M}(\rho(0),\rho(1))\geq 2\operatorname{BW}_{\partial M}(P_{b})>c,$
by Assumption 3. Let $J$ be the interval $[0,d_{\partial M}(\rho(0),\rho(1))]$. The path $\sigma$ has image in a subset of $C_{f}(N)$ isometric to
@@
Since $f$ is convex, the metric in $W$ dominates the metric in the Euclidean cone
@@
Also note that $W^{\prime}$ is isometric to the standard Euclidean cone
@@
Note that $|J|>c$, so the interval $\delta J$ has length greater than $\pi$. In particular the geodesic in $W^{\prime\prime}$ from $(\tau(0)-t_{0},0)$ to $(\tau(1)-t_{0},\delta|J|)$ goes through the cone point. Thus
$d_{M}(z,w)\geq(\tau(0)-t_{0})+(\tau(1)-t_{0}).$
There is a geodesic in $C_{f}(N)$ of length exactly $(\tau(0)-t_{0})+(\tau(1)-t_{0})$, but it goes through the cone point, contradicting our assumption that it doesn’t. This contradiction allows us to conclude that $\rho(s)\in P_{b}$, as desired.

### v2

@@ 5. Proof of Theorem B
We first claim that $B_{\varepsilon}(y)$ is convex in $M$. For $z,w\in B_{\varepsilon}(y)$ we must show any geodesic $[z,w]$ lies in $B_{\varepsilon}(y)$. By the triangle inequality, such a geodesic stays inside $B_{3\varepsilon}(y)$. Since $\varepsilon<\frac{1}{3}(b^{\prime}-t(y))$, we have $t(x)<b^{\prime}<\operatorname{BW}_{M}(\partial M)$ for any $x\in B_{3\varepsilon}(y)$. In particular the ball $B_{3\varepsilon}(y)$ is contained in $C_{f}(N)$ and the geodesic $[z,w]$ is also a geodesic in the warped product metric on $C_{f}(N)$ (as opposed to a path leaving $C_{f}(N)$). By Lemma 4.1, $C_{f}(N)$ is $\operatorname{CAT}(0)$. In particular, balls in $C_{f}(N)$ are convex, so the segment $[z,w]$ lies inside $B_{\varepsilon}(y)$ as desired.
The above argument also shows that restriction of $d_{X}$ to $B_{\varepsilon}(y)$ is equal to the restriction of $d_{C_{f}(N)}$. Since $f$ is a strictly increasing convex function, all the conclusions of Lemma 2.3 apply to any geodesic in $B_{\varepsilon}(y)$.
We now assume $z,w$ are in $Y\cap B_{\varepsilon}(y)$. Let $u$ be a point on $[z,w]$. We must show $u\in Y$. Using Lemma 2.3.(2) and our assumption on $\varepsilon$, we have
$t(u)\leq\max\{t(z),t(w)\}<b^{\prime}.$
To show $u\in Y$, we must show either that $p(u)\in P_{t(u)}$ (if $t(u)>b$), or that $p(u)\in P_{b}$ (if $t(u)\leq b$). To simplify, we introduce the following notation
@@
We have $P_{t(u)}\supseteq P_{t(z)}\cup P_{t(w)}$ by hypothesis 2, and so $P^{*}_{t(u)}\supseteq P^{*}_{t(z)}\cup P^{*}_{t(w)}$. Notably this means that the endpoints of the NN–geodesic $[p(z),p(w)]$ are in $P^{*}_{t(u)}$. Since $z$ and $w$ are in $B_{\varepsilon}(y)$,
@@
Because $\varepsilon<\frac{1}{2}L=\frac{1}{3}(t(y)-t_{0})$, the ball $B_{\varepsilon}(y)$ is disjoint from $t_{0},L]\times N\subset C_{f}(N)$. Since the warping function $f$ is increasing and $B_{\varepsilon}(y)$ is entirely above $\{L\}\times N$,
($\spadesuit$) $d_{X}(z,w)\geq d_{\{L\}\times N}((L,p(z)),(L,p(w)))=f(L)d_{N}(p(z),p(w)).$
@@
Without loss of generality suppose $t(w)<t(u)<t(z)$. By the argument in Case 1, the NN–geodesic $p(z),p(w)]\subseteq P^{*}_{t(w)}$. For the sake of contradiction, suppose that $u\in[z,w]$ is not in $Y$. Applying Case 1 and the convexity of the function $t(\cdot)$ (Lemma 2.3.(2)), we may re-choose $w$ so that $t(\cdot)$ is monotone on $[z,w]$. Possibly shortening the segment again, we may assume that $[z,w]$ meets $Y$ only in its endpoints. Thus, for $t\in(t(w),t(z))$, we must have ${P}^{*}_{t}\neq P^{*}_{t(w)}$. Hence $t(w)\geq b$, since $P^{*}_{t}=P_{b}$ for all $t\leq b$. This means that $[z,w]$ is actually a shortest path in $M$, contained in some $2\varepsilon$–ball around a point in $S$. Similarly, the points $z,w$ are points of $S$. Since $\varepsilon<\frac{1}{2}\mathrm{lcr}_{S}(S)$, all of $[z,w]$ is contained in $S$, and hence in $Y$. In particular, $u\in Y$, a contradiction. ∎
@@
Let $y$ be the cone point of $C_{f}(N)$ for a component $N$ of $\partial M$ so that $N\cap S\neq\emptyset$. Let $\varepsilon=\frac{1}{3}(b-t_{0}).$ For any $z,w\in B_{\varepsilon}(y)$, the triangle inequality shows that any geodesic $[z,w]$ lies entirely in $C_{f}(N)$, and
@@
Lemma 4.1 shows $C_{f}(N)$ is $\operatorname{CAT}(0)$, so in fact there is a unique geodesic $[z,w]$, which lies in $B_{\varepsilon}(y)$.
We suppose now that $z,w$ lie in $Y\cap B_{\varepsilon}(y)$. Note $Y\cap B_{\varepsilon}(y)$ consists of exactly those points $x$ so that $t(x)\leq t_{0}+\epsilon$ and $p(x)\in P_{b}$.
If $[z,w]$ goes through the cone point it is locally constant in the $N$ coordinate away from the cone point, by Lemma 2.3.(1), so it is clearly contained in $Y$.
Suppose then that $[z,w]$ avoids the cone point. We can write $[z,w]$ as a path
@@
where $\tau(s)>t_{0}$ for all $s$. The path $\rho(s)$ is a (not necessarily constant speed) geodesic in $N$ (see Lemma 2.3.(3)). We must show that $\rho(s)\in P_{b}$ for all $s$. By way of contradiction suppose some $\rho(s)\not\in P_{b}$. Replacing $[z,w]$ by a sub-segment if necessary, we may assume that $\rho(s)\not\in P_{b}$ for all $s\in(0,1)$. We therefore have
$d_{N}(\rho(0),\rho(1))\geq 2\operatorname{BW}_{N}(P_{b})>c,$
by Assumption 3. Let $J$ be the interval $0,d_{N}(\rho(0),\rho(1))]$. The path $\sigma$ has image in a subset of $C_{f}(N)$ isometric to
@@
Since $f$ is convex, the metric in $W$ dominates the metric in the warped product
@@
Also note that $W^{\prime}$ is isometric to the Euclidean cone on a rescaled copy of $J$,
@@
Note that $|J|>c$, so the interval $\delta J$ has length greater than $\pi$. This implies that the geodesic in $W^{\prime\prime}$ from $(\tau(0)-t_{0},0)$ to $(\tau(1)-t_{0},\delta|J|)$ goes through the cone point (see Lemma 2.3.(4)). Thus
$d_{X}(z,w)\geq(\tau(0)-t_{0})+(\tau(1)-t_{0}).$
But this implies that $d_{X}(z,w)=(\tau(0)-t_{0})+(\tau(1)-t_{0})$, since there is a path of exactly that length going through the cone point. (This path is locally constant in $N$ away from the cone point.) By uniqueness of geodesics in a $\operatorname{CAT}(0)$ space, this contradicts the assumption that $[z,w]$ misses the cone point.
This contradiction allows us to conclude that $\rho(s)\in P_{b}$, as desired. ∎
Appendix A The boundary at infinity of the coneoff.
In [Sw20], Świątkowski defines the tree of manifolds $\mathcal{X}(\mathcal{M})$ based on a collection of manifolds $\mathcal{M}$ as a certain inverse limit of connect sums of elements of $\mathcal{M}$. We refer to [Sw20] for the relevant definitions. We show in this appendix that the boundary at infinity of the universal cover of $\widehat{M}$ is such a tree of manifolds. Note that this boundary is also the Gromov boundary of $\pi_{1}\widehat{M}$.
Theorem A.1.
[…]
To prove Theorem A.1 we will need some more information about the cones described in Lemma 4.1. Recall that the space of directions at $p$, written $\Sigma_{p}$, is a set of equivalence classes of (nondegenerate) geodesic segments starting at $p$. Two geodesic segments are equivalent if the Alexandrov angle between them is zero. The Alexandrov angle makes $\Sigma_{p}$ into a metric space (see [BH, II.3] for more detail). At any point of a Riemannian manifold, the space of directions can be identified with the unit tangent sphere, and the Alexandrov angle is the angle given by the Riemannian metric at that point. The cone described in Lemma 4.1 is Riemannian everywhere except for the cone point, where the space of directions is described by the following lemma.
Lemma A.2.
[AB] Let $N$ and $C_{f}(N)$ be as in Lemma 4.1. The space of directions at the cone point is homeomorphic to $N$ and Alexandrov angle in this space is given by
$\angle(a,b)=\min\left\{\pi,\frac{1}{\delta}d_{N}(a,b)\right\}.$
Proof.
We recall that $f\colon\thinspace[t_{0},\infty)\to[0,\infty)$ is a cone warping function with parameters $(b,\delta,K)$, as in Definition 3.5. In particular the function $f$ vanishes at $t_{0}$, and $f^{\prime}(t_{0}^{+})=\delta$. Recall also that $C_{f}(N)$ is a warped product with fiber $N$ and warping function $f$. This description of the space of directions at the cone point of such a warped product follows from [FM, Proposition 3.7], which is a restatement of results from [AB, p. 1147]. ∎
Proof of Theorem A.1.
We must verify that the universal cover of our model $X\cong\widehat{M}$ is a Riemannian $(\mathcal{N},0)$–pseudomanifold with log-injective singularities, in the sense of [Sw20, Definition 6.1]. We can then apply [Sw20, Theorem 6.2] to deduce that the visual boundary is homeomorphic to the tree of manifolds $X(\mathcal{N})$ as desired. The definition [Sw20, Definition 6.1] consists of six conditions. The first four conditions are immediate from the fact that our space $X$ is a compact nonpositively curved pseudomanifold with isolated singularities, whose metric is Riemannian away from those singularities.
Świątkowski’s conditions (5) and (6) concern the space of directions $\Sigma_{p}$ at a point $p$ of $X$. Taking a small enough neighborhood $U$ of $p$ in $X$, there is a unique geodesic segment $\sigma_{x}$ from $p$ to any $x\in U$. This gives a well-defined logarithm
$\log_{p}(x)=([\sigma_{x}],d_{X}(p,x))$
mapping $U\smallsetminus p$ into $\Sigma_{p}\times(0,\infty)$.
Świątkowski’s Condition (5) is that each $p$ has a neighborhood $U$ so that $\log_{p}$ is injective on $U\smallsetminus\{p\}$. For $p$ non-singular, we may take $U$ to be $B_{\epsilon}(p)$, where $\epsilon$ is any number smaller than half the (Riemannian) injectivity radius at $p$; the logarithm is clearly injective on this neighborhood. For $p$ singular, we take $U$ to be a neighborhood of the form
$U\cong[t_{0},t_{1})\times_{f}N,$
where $t_{1}-t_{0}$ is small enough so that $U$ is convex. Now suppose two points $x=(t,n_{1})$ and $y=(s,n_{2})$ in $U\smallsetminus\{p\}$ satisfy $\log_{p}(x)=\log_{p}(y)$. Since the second factor of the logarithm records the distance, we have $t=s$. By Lemma 2.3.(1), each of the geodesics $\sigma_{x}$ and $\sigma_{y}$ has constant second coordinate. Lemma A.2 tells us that the Alexandrov angle between $\sigma_{x}$ and $\sigma_{y}$ is the maximum of $\pi$ and $\frac{1}{\delta}d_{N}(n_{1},n_{2})$. In particular it can only be zero if $n_{1}=n_{2}$ and $x=y$.
Świątkowski’s condition (6) is that, for any $p\in X$ and any $r\in(0,\pi)$, every ball of radius $r$ in $\Sigma_{p}$ with respect to the Alexandrov angle metric is a standard collared disk. This is immediate for $p$ nonsingular, since the space of directions in a Riemannian manifold is a standard round sphere. For $p$ equal to a cone point, we again use Lemma A.2. This implies that a ball of radius $r$ in $\Sigma_{p}$ can be identified with a ball of radius $\delta r$ in $N$. Since $\delta r$ is less than the injectivity radius of $N$, such a ball is indeed a collared disk. ∎